# Relative-error monotonicity testing

### v1

@@ 1 Introduction
i.e. the fraction of points in $\{0,1\}^{n}$ on which they disagree. This model is closely analogous to the standard “dense graph” property testing model [GGR98], in which the testing algorithm may make black-box queries for entries of the unknown graph’s adjacency matrix and the distance between two undirected $n$-node graphs is the fraction of all (${n\choose 2}$ possible edges that are present in one graph but not the other.
@@
What if we are interested in testing sparse Boolean functions, i.e. functions $f\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\to\{0,1\}$ that have $|f^{-1}(1)|\ll 2^{n}$? Analogous to the situation for graphs that was described above, every sparse Boolean function is Hamming-distance-close to the constant-0 function, so the standard Boolean function property testing model is not well suited for testing sparse functions.
@@
We observe that while the relative distance between two Boolean functions is not symmetric (i.e. $\mathsf{rel}\text{-}\mathsf{dist}(f,g)$ is not necessarily equal to $\mathsf{rel}\text{-}\mathsf{dist}(g,f)$), this does not pose a problem for us. The reason is that we are interested in the setting where $\mathsf{rel}\text{-}\mathsf{dist}(f,g)$ is small, and it is easily verified that if $\mathsf{rel}\text{-}\mathsf{dist}(f,g)=\varepsilon$ is at most (say) $0.99$, then $\mathsf{rel}\text{-}\mathsf{dist}(g,f)$ is also $O(\varepsilon)$; so in the regime of interest to us, relative distance is symmetric “up to constant factors.” We further observe that in the regime of our interest, the definition of relative distance is quite robust: as long as $\mathsf{rel}\text{-}\mathsf{dist}(f,g)$ is at most a small constant, then replacing $|f^{-1}(1)|$ in the denominator by any of $|g^{-1}(1)|$, $|f^{-1}(1)|+|g^{-1}(1)|$, $\max\{|f^{-1}(1)|,|g^{-1}(1)|\}$ or $\min\{|f^{-1}(1)|,|g^{-1}(1)|\}$ only changes the value of $\mathsf{rel}\text{-}\mathsf{dist}(f,g)$ by a small constant factor.
@@ 1.1 Our results
Upper and lower bounds for monotonicity testing: We focus on monotonicity testing in the relative-error model. Using the analysis of the “edge tester” [GGL+00] for monotonicity testing in the standard model, it is easy to obtain a one-sided [^2] Recall that a one-sided tester for a class of functions is one which must accept (with probability $1$) any function in the class. This is in contrast to making two-sided error, where an algorithm may reject a function in the class with small probability., non-adaptive [^3] A non-adaptive algorithm is one in which the choice of its $i$-th query point does not depend on the responses received to queries $11,\ldots,i-$ 1. algorithm with $O(n/\varepsilon)$ queries in the relative-error model (see Section 4.1 as a warmup). Our main algorithm as stated below, on the other hand, makes $O(\log(N)/\varepsilon)$ queries, where $N\mathrel{\mathop{\ordinarycolon}}=|f^{-1}(1)|$ denotes the sparsity of $f$ and is not given to the algorithm. Note that one always has $\log(N)/\varepsilon\leq n/\varepsilon$ as $N\leq 2^{n}$ trivially but the former can be asymptotically lower when $f$ is sparse (e.g., $\smash{n^{1/3}/\varepsilon}$ when $\smash{N=2^{n^{1/3}}}$).
@@
Summarizing our results, we see that there is both a correspondence and a significant point of difference between the state-of-the-art for standard-model monotonicity testing upper and lower bounds [CWX17, KMS18] and the upper and lower bounds for relative-error monotonicity testing that we establish. (We fix the error parameter $\varepsilon$ to be a constant, for simplicity, in the discussion below.) In both cases there is a 3/2-factor-gap between the exponent of the best known adaptive lower bound and the best known algorithm, but the source of this gap is intriguingly different between the two cases. In the standard model, the best known algorithm is the sophisticated “path tester” of Khot et al. [KMS18] which makes $\smash{\tilde{O}(\sqrt{n})}$ queries, while the strongest lower bound (for general testers, i.e. adaptive testers with two-sided error) is $\tilde{\Omega}(n^{1/3})$, due to Chen et al. [CWX17]. In contrast, in the relative-error setting the best algorithm we know of is a variant of the simple “edge tester” (given in Section 4), which achieves an $O(\log(N))$ complexity. But in the relative-error setting, for a wide range of values of $N=|f^{-1}(1)|$, it is possible to prove a stronger $\tilde{\Omega}((\log N)^{2/3})$ lower bound than in the standard model. So the gap in the exponent is again a factor of 3/2; but in the relative-error setting the state-of-the-art algorithm is a simple one (in contrast with the sophisticated algorithm and analysis, based on isoperimetry, from [KMS18]), and a stronger lower bound is achievable in the relative-error model than in the standard model.
@@ 1.2 Technical Overview
Now we consider the more challenging situation that the algorithm does not know the sparsity $N$. Our approach is to first obtain (an estimate of) the minimum Hamming weight of all points in $f^{-1}(1)$ by drawing a few samples from the sampling oracle, and then run the algorithm sketched above. In more detail, suppose that $x$ is the sampled point with minimum Hamming weight, which is $n-k$; our algorithm samples a few uniform random points from all points that are above $x$. If any of them have $f=0$ then the algorithm rejects, and otherwise it uses $k$ as a proxy for $\log N$ in the algorithm sketched above. To see why this works, it suffices to consider the case that $f$ is $\varepsilon$-relative-error-far from monotone. Observe first that if any sampled $x\in f^{-1}(1)$ has weight $\|x\|_{1}<n-2\log N$, then most points above $x$ cannot be in $f^{-1}(1)$ (since there are $N^{2}$ points above $x$ and only $N$ satisfying assignments in total), so sampling a few uniform random points above $x$ will reveal a violation of monotonicity. So we may assume that the sampled point $x$ with minimum Hamming weight has Hamming weight at least $n-2\log N$. Now our earlier analysis shows that the earlier algorithm will indeed reject any $f$ that has relative error at least $\varepsilon$ for monotonicity.
@@
Intuitively, to make monotonicity testing as hard as possible, we would like to use functions that have as little poset structure of the cube as possible. (For example, all monotonicity lower bounds in the standard model use functions that are only nontrivial in the middle layers, i.e., points with Hamming weight $(n/2)\pm O(\sqrt{n})$, with all points above the middle layers set to $1$ and all points below set to 00 by default. This way only the poset structure in the middle $O(\sqrt{n})$ layers is relevant.) Towards this end, our lower bound constructions use functions which are only nontrivial on two adjacent layers of the Boolean hypercube, we call them two-layer functions (see Section 5). More precisely, a function $f$ is a two-layer function if it is only nontrivial on points in layers $43n/$ 4 and $13n/4+$ 1: every point with weight $<3n/4$ is set to 00 and every point with weight $>3n/4+1$ is set to $1$. (We use the constant $3/4$ just to make the presentation more concrete; it can be replaced by any constant strictly between $1/2$ and $1$.)
In addition to the much simplified poset structure (which we discuss in the paragraph about our construction below), two advantages of using two-layer functions follow from simple calculations: (1) For $f$ to have $\varepsilon$-relative error from monotonicity for some constant $\varepsilon$, it suffices to show that a constant fraction of points in the two layers $43n/$ 4 and $13n/4+$ 1 must be changed to make the function monotone; (2) The use of two-layer functions let us reduce the analysis of general relative-error testing algorithms (which can use both $\mathrm{MQ}$ and $\mathrm{Samp}$ oracles) to algorithms that use $\mathrm{MQ}$ only (see 16).
So the main technical challenge is to establish strong lower bounds for two-layer functions. Recall that for monotonicity testing in the standard model, state of the art lower bounds are obtained using the Talagrand DNF [BB16, CWX17] (a random $2^{\sqrt{n}}$-term DNF in which each term has $\sqrt{n}$ variables sampled uniformly at random) and a depth-3 extension of the Talagrand DNF [CWX17]. The $\sqrt{n}$ (in both the size of each term and the exponent of the number of terms) comes from the fact we mentioned early about the middle $O(\sqrt{n})$ layers: if $\sqrt{n}$ were replaced by a larger polynomial $n^{0.5+c}$, then most layers in the middle would become trivial (points in layers above $(n/2)+O(n^{0.5-c})$ would be most likely set to $1$, and points in layers below $(n/2)-O(n^{0.5-c})$ would be most likely set to 00).
@@ 1.3 Related work
A similar model, in which the relative-distance measure is employed and the testing algorithm has both $MQ(f)$ and $\mathrm{Samp}(f)$ oracles, was considered by Ron and Tsur [RT14], who considered the problem of testing sparse images for properties such as sparsity, convexity, monotonicity, and being a line. In their setting, the domain of interest was the two-dimensional discrete grid $[n]\times[n]$, rather than the high-dimensional Boolean hypercube $\{0,1\}^{n}$ that we consider, and hence the results and techniques are very different between their setting and ours.
@@ 2 Preliminaries
Let $f\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\to\{0,1\}$ be the unknown Boolean function that is being tested. Recall that a sampling oracle for $f$ is an oracle which takes no inputs and, each time it is invoked, independently returns a uniform random element of $f^{-1}(1)$. [^5] If $f$ is the constant-0 function then the sampling oracle is assumed to return a special $\bot$ symbol; note that if this happens it completely identifies the function $f$ as the constant-0 function. Hence in our subsequent discussion we implicitly assume that the function $f$ being tested is not the constant-0 function. A black-box oracle for $f$ takes as input an $n$-bit string $x$ and returns $f(x)$; we sometimes refer to a black-box oracle simply “oracle for $f$.”
@@
We note that without loss of generality a relative-error testing algorithm can be assumed to make all $q$ of its calls to the sampling oracle before making any calls to the black-box oracle.
@@ 3 Some general results on relative-error testing
The query complexity of $T^{\prime}$ is clearly as claimed. To establish correctness, we begin by observing that by a simple Chernoff bound, with probability $99/100$ the value of $\bm{m}_{1}^{\prime}/m_{1}$ is within ±$\pm\varepsilon/100$ of the true value of $|f^{-1}(1)|/2^{n}$ (we will use this repeatedly in the following arguments).
@@
Algorithm $T^{\prime}$ first makes $m=O(1/p)$ calls to the black-box oracle for $f$ on uniform random inputs, and rejects if fewer than $43pm/$ 4 of the queried points are satisfying assignments of $f$. If $T^{\prime}$ does not reject in this first phase, it then performs $O(1)$ many runs of the standard-model testing algorithm $T$, each time with closeness parameter $p\varepsilon/2$, and outputs the majority of those runs.
@@
Let ${\cal C}$ be a class of functions from $\{0,1\}^{n}$ to $\{0,1\}$. Let $A$ be an algorithm which performs $\varepsilon$-relative-error proper learning of ${\cal C}$ using $s(\varepsilon,n)$ uniform samples from $f^{-1}(1).$ Then there is an $\varepsilon$-relative-error testing algorithm $T$ for ${\cal C}$ which makes at most $s(\varepsilon/4,n)+O(1/\varepsilon)$ calls to the sample oracle and at most $O(1/\varepsilon)$ calls to the black-box oracle $\mathrm{MQ}(f).$
@@
    Draw $m\mathrel{\mathop{\ordinarycolon}}=O(1/\varepsilon)$ samples from $\mathrm{Samp}(f)$ and evaluate $h$ on each of them. If $h$ evaluates to 0 on more than $(3/8)\varepsilon m$ of the samples, then output “reject.” Otherwise,
@@
Next, suppose that $\mathsf{rel}\text{-}\mathsf{dist}(f,{\cal C})>\varepsilon.$ If the relative-error learning algorithm does not output a hypothesis $h\in{\cal C}$, then the tester $T$ outputs “reject” (as desired), so suppose that in stage 1 the tester $T$ outputs a hypothesis $h\in{\cal C}$. Let $a,b,c\in[0,1]$ be as in Equation 1. Since ${\frac{b+c}{a+b}}=\mathsf{rel}\text{-}\mathsf{dist}(f,h)\geq\mathsf{rel}\text{-}\mathsf{dist}(f,{\cal C})>\varepsilon$, it must be the case that either ${\frac{b}{a+b}}$ (which is the fraction of examples in $f^{-1}(1)$ that are labeled 0 by $h$) is at least $135\varepsilon/$ 13, or ${\frac{c}{a+b}}$ is at least $138\varepsilon/$ 13. In the first case, a Chernoff bound gives that $T$ outputs “reject” in Step 2 with probability at least $2/3$, and in the second case,
@@
Note that $\frac{c}{a+c}$ is the fraction of examples in $h^{-1}(1)$ that are labeled 0 by $f$, so a Chernoff bound gives that $T$ outputs “reject” in Step 3 with probability at least $2/3$

### v2

@@ 1 Introduction
i.e. the fraction of points in $\{0,1\}^{n}$ on which they disagree. This model is closely analogous to the standard “dense graph” property testing model [GGR98], in which the testing algorithm may make black-box queries for entries of the unknown graph’s adjacency matrix and the distance between two undirected $n$-node graphs is the fraction of all $n\choose 2$} possible edges that are present in one graph but not the other.
@@
What if we are interested in testing sparse Boolean functions, i.e. functions $f\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\to\{0,1\}$ that have $f^{-1}(1)\ll 2^{n}$? Analogous to the situation for graphs that was described above, every sparse Boolean function is Hamming-distance-close to the constant-0 function, so the standard Boolean function property testing model is not well suited for testing sparse functions.
@@
We observe that while the relative distance between two Boolean functions is not symmetric (i.e. $\mathsf{rel}\text{-}\mathsf{dist}(f,g)$ is not necessarily equal to $\mathsf{rel}\text{-}\mathsf{dist}(g,f)$), this does not pose a problem for us. The reason is that we are interested in the setting where $\mathsf{rel}\text{-}\mathsf{dist}(f,g)$ is small, and it is easily verified that if $\mathsf{rel}\text{-}\mathsf{dist}(f,g)=\varepsilon$ is at most (say) $0.99$, then $\mathsf{rel}\text{-}\mathsf{dist}(g,f)$ is also $O(\varepsilon)$; so in the regime of interest to us, relative distance is symmetric “up to constant factors.” We further observe that in the regime of our interest, the definition of relative distance is quite robust: as long as $\mathsf{rel}\text{-}\mathsf{dist}(f,g)$ is at most a small constant, then replacing $|f^{-1}(1)|$ in the denominator by any of $|g^{-1}(1)|$, $|f^{1}(1)|+|g^{-1}(1)|$, $\max\{|f^{-1}(1)|,|g^{-1}(1)|\}$ or $\min\{|f^{-1}(1)|,|g^{-1}(1)|\}$ only changes the value of $\mathsf{rel}\text{-}\mathsf{dist}(f,g)$ by a small constant factor.
@@ 1.1 Our results
Upper and lower bounds for monotonicity testing: We focus on monotonicity testing in the relative-error model. Using the analysis of the “edge tester” [GGL+00] for monotonicity testing in the standard model, it is easy to obtain a one-sided [^2] Recall that a one-sided tester for a class of functions is one which must accept (with probability $1$) any function in the class. This is in contrast to making two-sided error, where an algorithm may reject a function in the class with small probability., non-adaptive [^3] A non-adaptive algorithm is one in which the choice of its $i$-th query point does not depend on the responses received to queries $1,\ldots,i-1$. algorithm with $O(n/\varepsilon)$ queries in the relative-error model (see Section 4.1 as a warmup). Our main algorithm as stated below, on the other hand, makes $O(\log(N)/\varepsilon)$ queries, where $N\mathrel{\mathop{\ordinarycolon}}=|f^{-1}(1)|$ denotes the sparsity of $f$ and is not given to the algorithm. Note that one always has $\log(N)/\varepsilon\leq n/\varepsilon$ as $N\leq 2^{n}$ trivially but the former can be asymptotically lower when $f$ is sparse (e.g., $\smash{n^{1/3}/\varepsilon}$ when $\smash{N=2^{n^{1/3}}}$).
@@
Summarizing our results, we see that there are both a correspondence and a significant point of difference between the state-of-the-art for standard-model monotonicity testing upper and lower bounds [CWX17, KMS18] and the upper and lower bounds for relative-error monotonicity testing that we establish. (We fix the error parameter $\varepsilon$ to be a constant, for simplicity, in the discussion below.) In both cases there is a 3/2-factor-gap between the exponent of the best known adaptive lower bound and the best known algorithm, but the source of this gap is intriguingly different between the two cases. In the standard model, the best known algorithm is the sophisticated “path tester” of Khot et al. [KMS18] which makes $\smash{\tilde{O}(\sqrt{n})}$ queries, while the strongest lower bound (for general testers, i.e. adaptive testers with two-sided error) is $\tilde{\Omega}(n^{1/3})$, due to Chen et al. [CWX17]. In contrast, in the relative-error setting the best algorithm we know of is a variant of the simple “edge tester” (given in Section 4), which achieves an $O(\log(N))$ complexity. But in the relative-error setting, for a wide range of values of $N=|f^{-1}(1)|$, it is possible to prove a stronger $\tilde{\Omega}((\log N)^{2/3})$ lower bound than in the standard model. So the gap in the exponent is again a factor of 3/2; but in the relative-error setting the state-of-the-art algorithm is a simple one (in contrast with the sophisticated algorithm and analysis, based on isoperimetry, from [KMS18]), and a stronger lower bound is achievable in the relative-error model than in the standard model.
@@ 1.2 Technical Overview
Now we consider the more challenging situation that the algorithm does not know the sparsity $N$. Our approach is to first obtain (an estimate of) the minimum Hamming weight of all points in $f^{-1}(1)$ by drawing a few samples from the sampling oracle, and then run the algorithm sketched above. In more detail, suppose that $x$ is the sampled point with minimum Hamming weight, which is nkn-k; our algorithm samples a few uniform random points from all points that are above $x$. If any of them have $f=0$ then the algorithm rejects, and otherwise it uses $k$ as a proxy for $\log N$ in the algorithm sketched above. To see why this works, it suffices to consider the case that $f$ is $\varepsilon$-relative-error-far from monotone. Observe first that if any sampled $x\in f^{-1}(1)$ has weight $\|x\|_{1}<n-2\log N$, then most points above $x$ cannot be in $f^{-1}(1)$ (since there are $N^{2}$ points above $x$ and only $N$ satisfying assignments in total), so sampling a few uniform random points above $x$ will reveal a violation of monotonicity. So we may assume that the sampled point $x$ with minimum Hamming weight has Hamming weight at least $n-2\log N$. Now our earlier analysis shows that the earlier algorithm will indeed reject any $f$ that has relative error at least $\varepsilon$ for monotonicity.
@@
Intuitively, to make monotonicity testing as hard as possible, we would like to use functions that have as little poset structure of the cube as possible. (For example, all monotonicity lower bounds in the standard model use functions that are only nontrivial in the middle layers, i.e., points with Hamming weight $(n/2)\pm O(\sqrt{n})$, with all points above the middle layers set to 11 and all points below set to 0 by default. This way only the poset structure in the middle $O(\sqrt{n})$ layers is relevant.) Towards this end, our lower bound constructions use functions which are only nontrivial on two adjacent layers of the Boolean hypercube, we call them two-layer functions (see Section 5.1). More precisely, a function $f$ is a two-layer function if it is only nontrivial on points in layers $3n/4$ and $3n/4+1$: every point with weight <3n/4<3n/4 is set to 0 and every point with weight >3n/4+1>3n/4+1 is set to $1$. (We use the constant $3/4$ just to make the presentation more concrete; it can be replaced by any constant strictly between $1/2$ and $1$).
In addition to the much simplified poset structure (which we discuss in the paragraph about our construction below), two advantages of using two-layer functions follow from simple calculations: (1) For $f$ to have $\varepsilon$-relative error from monotonicity for some constant $\varepsilon$, it suffices to show that a constant fraction of points in the two layers $3n/4$ and $3n/4+1$ must be changed to make the function monotone; (2) The use of two-layer functions helps us reduce the analysis of general relative-error testing algorithms (which can use both $\mathrm{MQ}$ and $\mathrm{Samp}$ oracles) to algorithms that use $\mathrm{MQ}$ only (see 15).
So the main technical challenge is to establish strong lower bounds for two-layer functions. Recall that for monotonicity testing in the standard model, state of the art lower bounds are obtained using the Talagrand DNF [BB16, CWX17] (a random $2^{\sqrt{n}}$-term DNF in which each term has $\sqrt{n}$ variables sampled uniformly at random) and a depth-3 extension of the Talagrand DNF [CWX17]. The $\sqrt{n}$ (in both the size of each term and the exponent of the number of terms) comes from the fact we mentioned early about the middle $O(\sqrt{n})$ layers: if $\sqrt{n}$ were replaced by a larger polynomial $n^{0.5+c}$, then most layers in the middle would become trivial (points in layers above $(n/2)+O(n^{0.5-c})$ would be most likely set to $1$, and points in layers below $(n/2)-O(n^{0.5-c})$ would be most likely set to 0).
@@ 1.3 Related work
A similar model, in which the relative-distance measure is employed and the testing algorithm has both $MQ(f)$ and $\mathrm{Samp}(f)$ oracles, was considered by Ron and Tsur [RT14], who considered the problem of testing sparse images for properties such as sparsity, convexity, monotonicity, and being a line. In their setting, the domain of interest was the two-dimensional discrete grid $n]\times[n]$, rather than the high-dimensional Boolean hypercube $\{0,1\}^{n}$ that we consider, and hence the results and techniques are very different between their setting and ours.
@@ 2 Preliminaries
Let $f\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\to\{0,1\}$ be the unknown Boolean function that is being tested. Recall that a sampling oracle for $f$ is an oracle which takes no inputs and, each time it is invoked, independently returns a uniform random element of $f^{-1}(1)$. [^5] If $f$ is the constant-0 function then the sampling oracle is assumed to return a special $\bot$ symbol; note that if this happens it completely identifies the function $f$ as the constant-0 function. Hence in our subsequent discussion we implicitly assume that the function $f$ being tested is not the constant-0 function. A black-box oracle for $f$ takes as input an $n$-bit string $x$ and returns $f(x)$; we sometimes refer to a black-box oracle as simple an “oracle for ff.”
@@
We note that without loss of generality a relative-error testing algorithm can be assumed to make all $q$ of its calls to the sampling oracle before making any calls to the black-box oracle. When we say that a relative-error testing algorithm is “non-adaptive”, this means that after receiving the results of all of its calls to the sampling oracle, it makes one parallel round of queries to the black-box oracle (so these queries can depend on the result of the calls to the sampling oracle, but the choice of the $i$-th query point for the black-box oracle does not depend on the responses received to queries $1,\ldots,i-1$).
@@ 3 Some general results on relative-error testing
The query complexity of $T^{\prime}$ is clearly as claimed. To establish correctness, we begin by observing that by a simple Chernoff bound, with probability $99/100$ the value of $\bm{m}_{1}^{\prime}/m_{1}$ is within $\pm\varepsilon/100$ of the true value of $|f^{-1}(1)|/2^{n}$ (we will use this repeatedly in the following arguments).
@@
Algorithm $T^{\prime}$ first makes $m=O(1/p)$ calls to the black-box oracle for $f$ on uniform random inputs, and rejects if fewer than $3pm/4$ of the queried points are satisfying assignments of $f$. If $T^{\prime}$ does not reject in this first phase, it then performs $O(1)$ many runs of the standard-model testing algorithm $T$, each time with closeness parameter $p\varepsilon/2$, and outputs the majority of those runs.
@@
Let $\cal C$} be a class of functions from $\{0,1\}^{n}$ to $\{0,1\}$. Let $A$ be an algorithm which performs $\varepsilon$-relative-error proper learning of $\cal C$} using $s(\varepsilon,n)$ uniform samples from $f^{-1}(1).$ Then there is an $\varepsilon$-relative-error testing algorithm $T$ for $\cal C$} which makes at most $s(\varepsilon/4,n)+O(1/\varepsilon^{2})$ calls to the sample oracle and at most $O(1/\varepsilon^{2})$ calls to the black-box oracle $\mathrm{MQ}(f).$
@@
    Draw $m\mathrel{\mathop{\ordinarycolon}}=O(1/\varepsilon^{2})$ samples from $\mathrm{Samp}(f)$ and evaluate $h$ on each of them. If $h$ evaluates to 0 on more than $(3/8)\varepsilon m$ of the samples, then output “reject.” Otherwise,
@@
Next, suppose that $\mathsf{rel}\text{-}\mathsf{dist}(f,{\cal C})>\varepsilon.$ If the relative-error learning algorithm does not output a hypothesis $h\in{\cal C}$, then the tester $T$ outputs “reject” (as desired), so suppose that in stage 1 the tester $T$ outputs a hypothesis $h\in{\cal C}$. Let $a,b,c\in[0,1]$ be as in Equation 1. Since ${\frac{b+c}{a+b}}=\mathsf{rel}\text{-}\mathsf{dist}(f,h)\geq\mathsf{rel}\text{-}\mathsf{dist}(f,{\cal C})>\varepsilon$, it must be the case that either $\frac{b}{a+b}$} (which is the fraction of examples in $f^{-1}(1)$ that are labeled 0 by $h$) is at least $5\varepsilon/13$, or $\frac{c}{a+b}$} is at least $8\varepsilon/13$. In the first case, a Chernoff bound gives that $T$ outputs “reject” in Step 2 with probability at least $2/3$, and in the second case,
@@
Note that $\frac{c}{a+c}$ is the fraction of examples in $h^{-1}(1)$ that are labeled 0 by $f$, so a Chernoff bound gives that $T$ outputs “reject” in Step 3 with probability at least $2/3.$